Chapter 17, Problem 047 A well with vertical sides and water at the bottom resonates at 8.40 Hz and at no lower frequency. (The air-filled portion of the well acts as a tuhe with one closed end and one open end.) The air in the well has a density of 1.10 kg/m3 and a bulk modulus of 1.33 × 105 Pa. How far down in the well is the water surface? Number the tolerance is t/2% Units
Answer #1

For a half closed pipe,

F = v/(4L)

Where, v = sqrt(B/p)

L = v/(4F)

= sqrt (B/p) /(4F)

= sqrt (1.33*10^5/1.10)/(4*8.4)

= 10.35 m